# IndicSuperTokenizer: An Optimized Tokenizer for Indic Multilingual LLMs

### v1

@@ 4.2 Extrinsic Evaluation on Downstream Tasks
Average 0.279 0.279 Average 0.388 0.394
4.3 How does tokenizer design impact model latency and throughput?
Next, we evaluate how tokenization impacts end-to-end model efficiency. We trained two 1B-parameter models under identical conditions: one with our tokenizer and one with the LLaMA tokenizer of similar vocabulary size. We then evaluated inference efficiency over 200 samples spanning Indic languages and English, with varying input lengths. Latency [^6] https://tinyurl.com/4e7nh7c8 was measured using standard metrics, including Time-To-First-Token (TTFT), Output Throughput (OTPT), and Input Sequence Length (ISL), across 200 instances with 5 warm-up requests and results averaged over 10 runs. Experiments were served on 8 H100 GPUs using Triton Inference Server as backend, with a maximum generation limit of 256 new tokens. Our tokenizer yields clear efficiency gains (Table 6). These gains stem from improved compression: shorter token sequences encode more information per token, thereby lowering per-request computation without compromising expressivity. Overall, this demonstrates that tokenizer design directly shapes not only pretraining efficiency but also real-world deployment latency, making it a critical factor for practical model performance.
@@ 4.3 How does tokenizer design impact model latency and throughput?
We allocate vocabulary budgets proportionally across scripts to preserve subword/multi-word granularity. Budgets were derived from corpus sizes, ensuring that both high- and low-resource scripts retained sufficient capacity. We compared two strategies for realizing this allocation. The first, explicit merging, trains script-specific tokenizers and concatenates their vocabularies via a rule-stacking procedure. While conceptually modular, this approach introduces distributional interference across scripts, yielding higher token-to-word ratios and fragmented segmentation (Table 9). The second, corpus-driven alignment, trains a single tokenizer on the concatenated multilingual corpus, allowing the vocabulary to adapt naturally to language frequencies. This unified training not only mirrored corpus composition (Table 10) but also achieved the lowest fertility scores across scripts (Table 3), outperforming explicit merging and public baselines. While script-aware budget allocation is necessary, explicit merging is inefficient; corpus-driven alignment provides a more scalable and faithful multilingual tokenization strategy.
Table 9: Fertility comparison between individual script tokenizers and the merged tokenizer across selected Indic languages. Lower values are better.
@@
Table 10: Script-specific training data size (Total corpus size 9.4 GB) and resulting vocabulary percentage distribution. Refer to Table 17 in Appendix for script mapping.
Metric ar bn deva en gu ka ml pa ta te
Data size (MB) 106 396 2200 3590 124 644 580 307 616 617
Percentage 1.12 4.18 23.25 37.94 1.31 6.81 6.13 3.24 6.51 6.52
Vocab perc dist 2.69 6.32 20.89 32.92 2.38 7.82 6.76 4.68 7.04 8.50
4.5 Quality Analysis: Undertrained “Glitch” Token
We analyze under-trained tokens in our tied-embedding LLaMA-3.2-1B models trained with both the IST tokenizer and a comparable BPE tokenizer of similar vocabulary size trained on the same corpus. Both tokenizers share the first 90% of the vocabulary. The IST tokenizer switches to superword training for the last 10% whereas the base BPE tokeniser continues standard subword training. Following land2024fishingmagikarpautomaticallydetecting to construct a reference for unused embeddings, we introduced a small set of dummy tokens into the vocabulary that have zero occurrences in the training data. Their embeddings were averaged to obtain a mean reference vector. We then retrieved the top-$K$ nearest neighbors (cosine distance), which represent potential “glitch” tokens (geiping2024coercing). As shown in Figure 5 (in the Appendix) the IST tokenizer produces far fewer such glitch tokens than the base BPE tokenizer. These results suggest that incorporating multi-words promotes more efficient utilization of the vocabulary, while purely subword-based tokenizers overfit in the long tail, yielding a higher proportion of under-trained tokens. More discussion in Appendix Section C.3.
4.6 Can we replace Opensource model tokenizer with IST?
Following ReTok (gu2024retokreplacingtokenizerenhance), we replace the tokenizer of a pre-trained LLaMA-3.2-1B model (denoted LLaMA-3.2-Orig(grattafiori2024llama) with IndicSuperTokenizer (referred as LLaMA-3.2-IST). Let $V_{\text{orig}}$ and $V_{\text{IST}}$ denote their corresponding vocabularies. For a token $t\in V_{\text{IST}}$, we initialize its embedding $E_{\text{init}}(t)$ as: if $t\in V_{\text{orig}}\cap V_{\text{IST}}$, then $E_{\text{init}}(t)=E_{\text{orig}}(t)$, its embedding from the pretrained model, otherwise, if $t\in V_{\text{IST}}\setminus V_{\text{orig}}$ and decomposes under the original tokenizer into $(t_{1},\ldots,t_{k})$, then $E_{\text{init}}(t)=\tfrac{1}{k}\sum_{i=1}^{k}E_{\text{orig}}(t_{i})$.
We then continually pretrained the LLaMA-3.2-IST model, keeping just the embedding and LM head layers trainable, on a 40B-token corpus comprising English, Indic, code, and mathematics (see Appendix for details). As seen in Table 11, the LLaMA-3.2-IST model performs competitively with the original LLaMA-3.2-ORIG. This suggests that, in addition to pretraining-from-scratch settings, an optimized multilingual tokenizer, such as IndicSuperTokenizer, could also be leveraged in opensource models through CPT (Continual Pretraining (chen2024towards)) leading to significant throughput gains (as seen in Table 6) while maintaining the original model quality.
Table 11: Performance comparison English (left) and Indic benchmarks (right).
English Benchmarks Indic Benchmarks
Dataset LLaMA-3.2-ORIG LLaMA-3.2-IST Dataset LLaMA-3.2-ORIG LLaMA-3.2-IST
Winogrande 0.60 0.61 Indic COPA 0.58 0.56
GSM8K 0.05 0.05 Indic Sentiment 0.82 0.85
ARC Challenge 0.40 0.39 Indic XNLI 0.35 0.34
MMLU 0.32 0.29 Indic Paraphrase 0.57 0.53
Average 0.34 0.34 Average 0.58 0.57
5 Ablation studies
Two-Stage vs. One-Stage: Controlling Vocabulary
Recently, BoundlessBPE (schmidt2024boundless) also explored a one-stage training paradigm in which pre-tokenization is governed by a fixed regular expression, enabling the direct learning of multiword units in a single pass. While effective in capturing frequent expressions, this strategy can also overfit to arbitrary character sequences lacking semantic value, ultimately reducing vocabulary efficiency. Our approach instead introduces a two-stage procedure. We replicate the one-stage setup of BoundlessBPE using its released regex (reffered as IST-BR) and compare against our two-stage tokenizer. As shown in Table 12, our method consistently achieves lower fertility across the top 10 Indic languages and English, indicating more compact and semantically grounded vocabularies. Overall, the comparison highlights a clear trade-off: while one-stage methods capture surface-level patterns indiscriminately, our two-stage design balances efficiency and linguistic integrity by decoupling subword and multiword learning.
Table 12: Fertility score ($\downarrow$) comparison between one-stage and two-stage IST tokenizers.
IST-BR (200K) 1.86 1.76 2.05 1.75 1.62 1.37 2.20 1.86 1.39 1.39 2.19 1.61 2.29 2.30 1.66 1.67 1.69 1.49 2.68 3.61 1.56 2.12 1.88 1.54
IST (180K/200K) 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Table 13: Ablation of tokenizer training data size and its impact on fertility score ($\downarrow$).
Size as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd Average
@@ 5 Ablation studies
Similar to (reddy2025much), we study the effects of scaling training data, however only in Stage 1 of our training. Figure 13 shows that our performance plateaus after 10G of data.
@@
We ablate the transition point $t$ (Section 3.1) at which training shifts from subword to cross-word merges. Varying $t$ reveals a clear trade-off: early transitions favor frequent multiword expressions but weaken morphological coverage, while late transitions preserve subwords at the cost of longer sequences. Across Indic and non-Indic languages, intermediate values of 90% $t$ yield the best balance, improving token efficiency and cross-lingual consistency (Table 14).
Table 14: Impact of varying transition point (as a $\%$ of vocab size 200200K) on fertility ($\downarrow$).
Transition (%) as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
60 1.91 1.80 2.05 1.39 1.38 1.04 2.16 1.77 1.16 1.15 2.17 1.53 2.30 2.29 1.56 1.58 1.68 1.39 2.48 3.89 1.43 2.11 1.86 1.45
75 1.91 1.79 2.05 1.41 1.38 1.04 2.16 1.77 1.16 1.15 2.16 1.53 2.30 2.28 1.56 1.58 1.68 1.39 2.47 3.91 1.43 2.10 1.86 1.45
80 1.89 1.78 2.03 1.41 1.38 1.05 2.15 1.77 1.16 1.16 2.14 1.53 2.28 2.26 1.56 1.57 1.67 1.39 2.46 3.83 1.42 2.08 1.83 1.44
85 1.87 1.77 2.01 1.43 1.39 1.06 2.13 1.76 1.17 1.16 2.13 1.53 2.26 2.25 1.56 1.56 1.66 1.39 2.46 3.78 1.42 2.07 1.82 1.44
90 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
95 1.85 1.75 1.98 1.47 1.42 1.10 2.13 1.74 1.21 1.20 2.12 1.53 2.23 2.24 1.56 1.56 1.66 1.41 2.46 3.68 1.43 2.06 1.81 1.44
@@
Vocabulary size strongly influences tokenization-model efficiency with trade-offs. Smaller vocabularies yield finer subword units that generalize well to unseen words but lengthen sequences, raising compute costs. Larger vocabularies shorten sequences by encoding frequent forms as single tokens, but waste capacity on rare items, inflate embeddings and softmax layers (shazeer2017outrageouslylargeneuralnetworks), and bias toward high-resource languages, hurting multilingual balance. With the same transition point at 90%, we found no significant impact on fertility scores beyond 200K (Table 15).
Table 15: Ablation on vocab size ($t=90\%$) and its impact on fertility ($\downarrow$) scores.
Vocab Size as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd
162K/180K 1.89 1.78 2.08 1.48 1.47 1.13 2.21 1.80 1.24 1.22 2.22 1.60 2.35 2.27 1.65 1.65 1.68 1.42 2.62 3.84 1.48 2.16 1.91 1.47
180K/200K 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.27 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
202K/225K 1.81 1.70 1.99 1.44 1.43 1.10 2.14 1.72 1.20 1.19 2.14 1.55 2.24 2.21 1.59 1.59 1.60 1.36 2.55 3.59 1.41 2.08 1.82 1.41
225K/250K 1.78 1.67 1.95 1.42 1.42 1.09 2.11 1.69 1.19 1.17 2.10 1.53 2.20 2.17 1.57 1.57 1.57 1.34 2.52 3.45 1.38 2.04 1.77 1.38
Effect of Normalization in Multilingual Tokenization
Unicode normalization is crucial for multilingual settings (karthika2025multilingualtokenizationlensindian), particularly for Indic languages, where a single grapheme can be represented by multiple Unicode sequences (e.g., pre-composed characters vs. base-plus-diacritic sequences), causing token fragmentation and inflated vocabulary size. Table 16 shows that NFKC yielded marginal but consistent gains by unifying character forms. Accordingly, we adopt NFKC to reduce variability and improve tokenizer robustness.

### v2

@@ 4.1 Intrinsic Evaluation of Tokenizers
Average 0.279 0.279
Indic Benchmarks
Indic COPA 0.544 0.556
Indic Sentiment 0.524 0.551
Indic XNLI 0.347 0.346
Indic Paraphrase 0.534 0.539
MILU (Indic Multi-turn LU) 0.261 0.258
ARC Challenge (Indic) 0.236 0.244
TriviaQA (Indic) 0.268 0.262
Average 0.388 0.394
Table 6: Extrinsic evaluation of MUTANT-Indic vs LLaMA-4 on English and Indic benchmarks.
4.2 Extrinsic Evaluation
We also evaluate the downstream model performance (see Table 6) by pretraining LLaMA-3.2 1B models using two tokenizers: i) MUTANT-Indic, our proposed tokenizer optimized for morphologically meaningful segmentation in Indic and multilingual settings, and (ii) LLaMA-4 tokenizer, chosen for comparable vocabulary size and widespread use. Both models were trained on the same dataset in iso-compute setting to ensure a fair comparison. More details in the Appendix B. We find that our tokenizer shows competitive performance across the English and Indic benchmarks. We additionally report Bits-per-Character in Appendix C.6, which remains meaningful even near random accuracy. We also trained a model using the Stage-1 tokenizer, which also attains strong downstream performance. Table 26 in Appendix shows that the Stage-1 tokenizer itself constitutes a strong and competitive baseline.
The pretraining corpus (Table 17 in Appendix) balances coverage and domain diversity. It combines web-scale sources (Nemotron CC) for general context with structured data including MegaMath (Zhou et al., 2025), StackV2 (Lozhkov et al., 2024), synthetic generations, and books. Indic-language content constitutes roughly $20\%$ of the corpus, drawn from Indic CC, Wikipedia, and Sangraha Verified (Khan et al., 2024), providing sufficient signal to evaluate cross-lingual and morphologically rich representation quality.
Gemma-3 GPT-OSS LLaMA-4 Sarvam Sutra MUTANT-Indic
Entropy $\downarrow$ 20.70 20.81 21.09 20.71 20.62 20.42
Efficiency $\uparrow$ 0.22 0.19 0.14 0.21 0.23 0.28
Model TTFT (ms) $\downarrow$ OTPT (tokens/s) $\uparrow$
LLaMA-4 19.17 $\pm$ 0.15 117.99
MUTANT-Indic 18.98 $\pm$ 0.36 169.42
Table 8: Inference latency comparison of 1B models trained with LLaMA-4 and MUTANT-Indic tokenizers.
Metric ar bn deva en gu ka ml pa ta te
Data size (MB) 106 396 2200 3590 124 644 580 307 616 617
Percentage 1.12 4.18 23.25 37.94 1.31 6.81 6.13 3.24 6.51 6.52
Vocab % 2.69 6.32 20.89 32.92 2.38 7.82 6.76 4.68 7.04 8.50
Table 9: Script-specific training data size (Total corpus size 9.4 GB) and resulting vocabulary % distribution.
4.3 Impact on latency and throughput
Next, we evaluate how tokenization impacts end-to-end model efficiency. We train two 1B-parameter models under identical conditions: one with our tokenizer and one with the LLaMA tokenizer of similar vocabulary size. We then evaluate inference efficiency over 200 samples spanning Indic languages and English, with varying input lengths. Latency [^5] https://tinyurl.com/4e7nh7c8 was measured using standard metrics, including Time-To-First-Token (TTFT), Output Throughput (OTPT), and Input Sequence Length (ISL), across 200 instances ( See Appendix C.4 for details) with 5 warm-up requests and results averaged over 10 runs. Experiments were done on 8 H100 GPUs using Triton Inference Server as backend, with a maximum generation limit of 256 new tokens. Our tokenizer yields clear efficiency gains (Table 8) which stem from improved compression: shorter token sequences encode more information per token, thereby lowering per-request computation without compromising expressivity. Overall, this demonstrates that tokenizer design directly shapes not only pretraining efficiency but also real-world deployment latency, making it a critical factor for practical model performance.
@@ 4.3 Impact on latency and throughput
We compare explicit vocabulary merging with corpus-driven joint training under script-aware budget constraints. Explicitly merging script-specific tokenizers leads to fragmented segmentation and higher fertility due to cross-script interference (Table 10). In contrast, corpus-driven joint training naturally aligns vocabulary allocation with data distribution (Table 9), achieving lower fertility and outperforming merged and public baselines (Table 3). These results indicate that joint, corpus-driven training is a more effective and scalable strategy.
@@
MUTANT-Indic 1.85 1.74 1.23 1.58 1.63 2.59 1.88
Table 10: Fertility comparison ($\downarrow$) between individual script tokenizers and the merged tokenizer across selected Indic languages. Lower values are better.
Dataset w/ Original LLaMA-4 w/ MUTANT-Indic
English Benchmarks
Winogrande 0.60 0.61
GSM8K 0.05 0.05
ARC Challenge 0.40 0.39
MMLU 0.32 0.29
Average 0.34 0.34
Indic Benchmarks
Indic COPA 0.58 0.56
Indic Sentiment 0.82 0.85
Indic XNLI 0.35 0.34
Indic Paraphrase 0.57 0.53
Average 0.58 0.57
Table 11: Performance comparison on English and Indic benchmarks in Continual Pretraining (CPT) setting of LLaMA-3.2 1B model (best scores in bold).
Tokenizer as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd Avg
Single-stage (200K) 1.86 1.76 2.05 1.75 1.62 1.37 2.20 1.86 1.39 1.39 2.19 1.61 2.29 2.30 1.66 1.67 1.69 1.49 2.68 3.61 1.56 2.12 1.88 1.54 1.90
Two-stage (180K/200K) 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44 1.83
Table 12: Fertility score ($\downarrow$) for single vs two-stage curriculum training (same evaluation data as Section 3.4).
Parameter as bn brx code doi eng gom gu hi kas kn mai ml mni mr nep or pa san sat snd ta te urd Average
Data size
@@
Transition point
60 1.91 1.80 2.05 1.39 1.38 1.04 2.16 1.77 1.16 1.15 2.17 1.53 2.30 2.29 1.56 1.58 1.68 1.39 2.48 3.89 1.43 2.11 1.86 1.45 1.81
75 1.91 1.79 2.05 1.41 1.38 1.04 2.16 1.77 1.16 1.15 2.16 1.53 2.30 2.28 1.56 1.58 1.68 1.39 2.47 3.91 1.43 2.10 1.86 1.45 1.81
80 1.89 1.78 2.03 1.41 1.38 1.05 2.15 1.77 1.16 1.16 2.14 1.53 2.28 2.26 1.56 1.57 1.67 1.39 2.46 3.83 1.42 2.08 1.83 1.44 1.80
85 1.87 1.77 2.01 1.43 1.39 1.06 2.13 1.76 1.17 1.16 2.13 1.53 2.26 2.25 1.56 1.56 1.66 1.39 2.46 3.78 1.42 2.07 1.82 1.44 1.80
90 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44 1.83
95 1.85 1.75 1.98 1.47 1.42 1.10 2.13 1.74 1.21 1.20 2.12 1.53 2.23 2.24 1.56 1.56 1.66 1.41 2.46 3.68 1.43 2.06 1.81 1.44 1.79
Vocab size
162K/180K 1.89 1.78 2.08 1.48 1.47 1.13 2.21 1.80 1.24 1.22 2.22 1.60 2.35 2.27 1.65 1.65 1.68 1.42 2.62 3.84 1.48 2.16 1.91 1.47 1.86
180K/200K 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.27 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44 1.82
202K/225K 1.81 1.70 1.99 1.44 1.43 1.10 2.14 1.72 1.20 1.19 2.14 1.55 2.24 2.21 1.59 1.59 1.60 1.36 2.55 3.59 1.41 2.08 1.82 1.41 1.79
225K/250K 1.78 1.67 1.95 1.42 1.42 1.09 2.11 1.69 1.19 1.17 2.10 1.53 2.20 2.17 1.57 1.57 1.57 1.34 2.52 3.45 1.38 2.04 1.77 1.38 1.75
Table 13: Ablation of tokenizer training data size (in GB), transition point (as a $\%$ of vocab size 200200K), vocab size ($t=90\%$) and its impact on fertility score ($\downarrow$). We use the same evaluation data as described in Section 3.4.
4.5 Quality Analysis: Undertrained Tokens
We analyze under-trained “Glitch” tokens in our tied-embedding LLaMA-3.2-1B models trained with both the MUTANT-Indic tokenizer and a comparable BPE tokenizer of similar vocabulary size trained on the same corpus. Both tokenizers share the first 90% of the vocabulary. The MUTANT-Indic tokenizer switches to multi-word training for the last 10% whereas the base BPE tokeniser continues standard subword training. Following Land and Bartolo (2024) to construct a reference for unused embeddings, we introduced a small set of dummy tokens into the vocabulary that have zero occurrences in the training data. Their embeddings were averaged to obtain a mean reference vector. We then retrieve the top-$K$ nearest neighbors (cosine distance), which represent potential “glitch” tokens (Geiping et al., 2024). As shown in Figure 6 (in the Appendix C.3), the MUTANT-Indic tokenizer produces far fewer such glitch tokens than the base BPE tokenizer. These results suggest that incorporating multi-words promotes more efficient utilization of the vocabulary, while purely subword-based tokenizers overfit in the long tail, yielding a higher proportion of under-trained tokens.
4.6 Can we replace Opensource model tokenizer with MUTANT-Indic?
Following ReTok (Gu et al., 2024), we replace the tokenizer of a pre-trained LLaMA-3.2-1B model (denoted LLaMA-3.2-Orig(Grattafiori et al., 2024) with MUTANT-Indic (referred as LLaMA-3.2-MUTANT-Indic). Let $V_{\text{orig}}$ and $V_{\text{MUTANT-Indic}}$ denote their corresponding vocabularies. For a token $t\in V_{\text{MUTANT-Indic}}$, we initialize its embedding $E_{\text{init}}(t)$ as: if $t\in V_{\text{orig}}\cap V_{\text{MUTANT-Indic}}$, then $E_{\text{init}}(t)=E_{\text{orig}}(t)$, its embedding from the pretrained model, otherwise, if $t\in V_{\text{MUTANT-Indic}}\setminus V_{\text{orig}}$ and decomposes under the original tokenizer into $(t_{1},\ldots,t_{k})$, then $E_{\text{init}}(t)=\tfrac{1}{k}\sum_{i=1}^{k}E_{\text{orig}}(t_{i})$.
We then continually pretrained the LLaMA-3.2-MUTANT-Indic model, keeping just the embedding and LM head layers trainable, on a 40B-token corpus comprising English, Indic, code, and mathematics (see Appendix for details). As seen in Table 11, the LLaMA-3.2-MUTANT-Indic model performs competitively with the original LLaMA-3.2-ORIG. This suggests that, in addition to pretraining-from-scratch settings, an optimized multilingual tokenizer, such as MUTANT-Indic, could also be leveraged in opensource models through CPT (Continual Pretraining (Chen et al., 2024)) leading to significant throughput gains (see Table 8).
5 Ablation studies
Two-Stage vs. One-Stage: Controlling Vocabulary
Recently, BoundlessBPE (Schmidt et al., 2024) also explored a one-stage training paradigm in which pre-tokenization is governed by a fixed regular expression, enabling the direct learning of multiword units in a single pass for English language. Our approach instead introduces a two-stage procedure. We replicate the one-stage setup of BoundlessBPE using its released regex and compare against our strategy. Our method consistently achieves lower fertility across the top 10 Indic languages and English (Table 12). Overall, the comparison highlights a clear trade-off: while one-stage methods capture surface-level patterns indiscriminately, our two-stage design balances efficiency and linguistic integrity by decoupling subword and multiword learning.
@@ 5 Ablation studies
Similar to Reddy et al. (2025), we study the effects of scaling training data in Stage 1 of our training. Table 13 shows that the performance plateaus after 10G of data.
@@
We ablate the transition point $t$ (Section 3.2) at which training shifts from subword to cross-word merges. Varying $t$ reveals a clear trade-off: early transitions favor frequent multiword expressions but weaken morphological coverage, while late transitions preserve subwords at the cost of longer sequences. Across Indic and non-Indic languages, intermediate values of 85-90% $t$ yields the best balance, improving token efficiency and cross-lingual consistency (Table 13).
@@
Vocabulary size strongly influences tokenization-model efficiency with trade-offs. Smaller vocabularies yield finer subword units that generalize well to unseen words but with increased compute costs. Larger vocabularies shorten sequences by encoding frequent forms as single tokens, but waste capacity on rare items, inflate embeddings and softmax layers (Shazeer et al., 2017), and have bias toward high-resource languages, hurting multilingual balance. With the same transition point at 90%, we found no significant impact on fertility scores beyond 200K (Table 13).
Effect of Normalization
Prior works show that unicode normalization is crucial for multilingual settings (Karthika et al., 2025b), which reduces token fragmentation and inflated vocabulary size. Table 25 (in Appendix) shows that NFKC yielded marginal but consistent gains by unifying character forms. Accordingly, we adopt NFKC to reduce variability and improve tokenizer robustness.